OpenStudy (anonymous):

solve each system by substitution or elimination. 1. y=2x+8 y=3x-1 2. 2x-y=2 2x-2y=4 3. -x+y=2 2x+y=-1

OpenStudy (anonymous):

Never put your question like that, it confuses the person helping you. (1).x=20 y=48 (2).y=-41 x=-39/2 (3).y=1 x=-1